# Learning WENO for entropy stable schemes to solve conservation laws

### v1

@@ Abstract
Entropy conditions play a crucial role in the extraction of a physically relevant solution for a system of conservation laws, thus motivating the construction of entropy stable schemes that satisfy a discrete analogue of such conditions. TeCNO schemes (Fjordholm et al. 2012) [14] form a class of arbitrary high-order entropy stable finite difference solvers, which require specialized reconstruction algorithms satisfying the sign property at each cell interface. Recently, third-order WENO schemes called SP-WENO (Fjordholm and Ray, 2016) [16] and SP-WENOc (Ray, 2018) [44] have been designed to satisfy the sign property. However, these WENO algorithms can perform poorly near shocks, with the numerical solutions exhibiting large spurious oscillations. In the present work, we propose a variant of the SP-WENO, termed as Deep Sign-Preserving WENO (DSP-WENO), where a neural network is trained to learn the WENO weighting strategy. The sign property and third-order accuracy are strongly imposed in the algorithm, which constrains the WENO weight selection region to a convex polygon. Thereafter, a neural network is trained to select the WENO weights from this convex region with the goal of improving the shock-capturing capabilities without sacrificing the rate of convergence in smooth regions. The proposed synergistic approach retains the mathematical framework of the TeCNO scheme while integrating deep learning to remedy the computational issues of the WENO-based reconstruction. We present several numerical experiments to demonstrate the significant improvement with DSP-WENO over the existing variants of WENO satisfying the sign property.
@@ 1 Introduction
Hyperbolic systems of conservation laws dictate the behavior of quantities that are conserved in time as the system evolves. These systems of PDEs are ubiquitous across many disciplines. A generic system of conservation laws in one spatial dimension can be mathematically expressed as
@@
However, it is well-known that solutions to nonlinear conservation laws can develop discontinuities in finite time even with smooth initial conditions. Thus, solutions to (1) must be understood in the weak (distributional) sense. Moreover, since weak solutions are not unique in general, additional constraints in terms of the entropy conditions must be imposed such that a physically relevant weak solution is extracted. Assume that for (1), $\eta({\bm{u}})$ is a convex entropy function and $q({\bm{u}})$ is the entropy flux function satisfying the compatibility condition
$\nabla_{{\bm{u}}}q({\bm{u}})^{\top}=\nabla_{{\bm{u}}}\eta({\bm{u}})^{\top}\nabla_{\bm{u}}{\bm{f}}({\bm{u}}).$
The solution ${\bm{u}}$ of (1) is said to be an entropy solution if it satisfies the following inequality for all admissible entropy pairs $(\eta({\bm{u}}),q({\bm{u}}))$ associated with (1):
@@
which is understood in the weak sense. The entropy is conserved for smooth solutions, and thus (2) is taken to be an equality, i.e., $\eta$ is the solution of a conservation law. In contrast, entropy is dissipated near shocks, thus the inequality in (2) is strict near such discontinuities. Scalar conservation laws have unique entropy solutions [27], but the uniqueness of entropy solutions is not guaranteed for general systems of conservation laws [31, 8]. However, the entropy conditions provide the only non-linear estimates currently available for generic systems of conservation laws [10], and are thus essential. In practice, the entropy condition (2) is considered with respect to a specific choice of $(\eta({\bm{u}}),q({\bm{u}}))$ for a general system (1).
@@
It is meaningful to develop entropy stables scheme, i.e., schemes satisfying a discrete version of (2), to solve systems of conservation laws. Further, it is essential for such methods to be high-order accurate while being capable of capturing discontinuities without spurious oscillations. There has been much focus on this area in the past few decades, and finite difference methods remain one such popular framework. In finite difference methods, the computational domain is partitioned into cells, on which a discrete version of the conservation law is formulated. Point values of the solution in the cells are evolved in time using time integration techniques, such as strong stability preserving Runge-Kutta methods [18]. In [53], a novel approach to constructing entropy stable finite difference schemes was proposed, which comprises two steps: (1) beginning with a second-order entropy conservative flux such that entropy is conserved locally and (2) adding an artificial dissipation term to ensure entropy stability.
TeCNO schemes [14] are a popular class of arbitrary high-order entropy stable finite difference schemes which build on the formulation considered in [53]. These schemes consist of augmenting high-order entropy conservative fluxes with high-order numerical diffusion, which rely on specialized polynomial reconstruction algorithms. Crucially, these reconstructions must satisfy a sign property at each cell interface so that entropy stability is ensured. This property requires the jump in the reconstructed cell interface values to have the same sign as the jump in the corresponding cell values.
@@
Although any reconstruction method with the sign property can be used to formulate entropy stable TeCNO schemes, only a handful of such algorithms are currently available. The second-order limited total variation diminishing (TVD) using the minmod limiter is the simplest reconstruction possessing the sign property [14]. Third-order and fourth-order sign-preserving reconstructions based on non-linear limiting of polynomials were proposed [7, 6]. The essentially non-oscillatory (ENO) reconstruction method, which adaptively chooses the smoothest stencil for reconstruction, satisfies the sign property as demonstrated in [15]. To overcome some of the computational challenges encountered when using ENO, third-order weighted ENO (WENO) reconstructions were proposed in [16, 44] which guarantee the sign property. However, these WENO schemes (termed as SP-WENO), while ensuring entropy stability in the TeCNO framework, are susceptible to spurious oscillations near discontinuities. We will highlight this issue in greater detail in the present work. We also mention here an alternate strategy [39] based on TeCNO schemes, where the sign property of reconstructions is not required. Instead, a specialized diffusion operator is constructed to ensure entropy stability, which can be used with any high-order reconstruction algorithm. We will not focus on this approach in the present work.
@@
The last few years have witnessed a surge in the use of deep learning-based strategies to solve problems in scientific computing. Key examples includes building surrogates to solve PDEs [43, 9, 35, 33, 40], learning parameter-to-observable maps [37, 36, 21], solving Bayesian inference problems [17, 41, 56, 11, 47], or learning closure models [51, 2, 38]. Among these is the class of deep learning (DL) approaches which aim to enhance existing numerical solvers, instead of replacing them. The philosophy here is to first identify computational bottlenecks in the solver, and then use domain knowledge to train specialized neural networks to replace the bottleneck, keeping the rest of the solver intact. Thus, this synergistic approach leverages the “best of both worlds”, leading to a more robust numerical solver aided by DL. For instance, in [45, 46], deep-learned troubled-cell indicators were designed for discontinuous Galerkin (DG) schemes solving conservation laws, to identify (classify) elements containing discontinuities. Once flagged, the DG solver can apply any suitable limiter in the troubled-cells to avoid Gibbs oscillations. These DL indicators are free of problem-dependent parameters and trained in model agnostic manner, i.e., the same network can be used to solve any system of conservation laws. Similar DL-based shock capturing strategies have also been explored in [13, 3, 58, 49, 4, 1]
@@
DL techniques have also been used to learn ENO/WENO-type reconstructions. In [12], it was shown that ENO (and some of its variants) are equivalent to deep ReLU neural networks, which in turn highlighted the potential of deep networks to approximate rough functions. A six-point ENO-type reconstruction was proposed in [32] where the stencil selection is performed using a neural network. However, when used to solve conservation laws, some numerical experiments were accompanied by a degeneracy in the expected order of accuracy. In [52], a neural network was trained to develop an optimal variant of the classical WENO-JS [23]. The trained network, called WENO-NN, was shown to outperform WENO-JS by being for anti-diffusive leading to sharper shock profiles in numerical solutions. However, this also lead to the development of spurious oscillations in under-resolved simulations. Further, WENO-NN was not able to retain the desired order of accuracy for smooth solutions. A DL-based fifth-order WENO called WENO-DS was proposed in [26], where the WENO weights were fine-tuned to improve its shock-capturing capabilities. The authors proved the formal order of accuracy of WENO-DS, and also demonstrated it numerically. However, WENO-DS is trained using the flux values as input to the network, thus it is designed to work for a specific conservation law model. The network is tuned separately for each PDE considered, and thus, one instance of WENO-DS is not generally model agnostic.
@@
In the present work, we are interested in using DL to recover reconstruction algorithms constrained to satisfy useful physical properties. In particular, we construct a variant of SP-WENO, which we call DSP-WENO, where a neural network predicts the weights associated with the reconstruction, while ensuring that the sign property (along with other crucial constraints) are satisfied. These constraints are strongly imposed which results in a convex polygonal region for the WENO weight selection. Then, a neural network is trained on suitably generated training samples to learn the weight selection algorithm of the DSP-WENO for appropriate behavior near discontinuities and smooth regions. The proposed DSP-WENO is constructed to be third-order accurate, which is also demonstrated numerically. When used in the TeCNO framework, DSP-WENO overcomes the computational issues faced by the existing SP-WENO strategies. We re-iterate that the DSP-WENO only replaces the reconstruction needed in the high-order diffusion term, keeping the rest of the TeCNO framework intact. Thus the proposed method can be seen as a deep learning-based enhancement of an existing numerical algorithm. We remark here that a single DSP-WENO network is trained (offline) to be used with any conservation laws, i.e., the proposed algorithm is agnostic to the specific PDE model being solved.
@@ 2 Finite Difference Schemes/Entropy Stable Schemes
We consider a one-dimensional finite difference formulation for ease of discussion, which can easily be extended to higher dimensions. We partition the spatial domain $\Omega$ using disjoint cells $I_{i}=[x_{i-\frac{1}{2}},x_{i+\frac{1}{2}})$ of uniform length $h\equiv x_{i+\frac{1}{2}}-x_{i-\frac{1}{2}}$. The mesh point $x_{i}$ is the center of the cell $I_{i}$ while the half indices denote the cell interfaces. Figure 1 shows an example of the discretized mesh. Keeping time continuous, the semi-discrete finite difference scheme for (1) is expressed as
@@
Here, ${\bm{u}}_{i}(t)$ approximates the point values of the solution to (1) at cell center $x_{i}$ and time $t$, while ${\bm{f}}_{i+\frac{1}{2}}$ is a consistent, conservative numerical approximation of the flux ${\bm{f}}$ at the cell interface $x_{i+\frac{1}{2}}$. We seek entropy stable schemes to approximate (1) which satisfy a discrete version of the entropy condition (2). As described in [53], we begin by constructing an entropy conservative scheme.
Definition 1.
The numerical scheme (3) is entropy conservative if it satisfies the following discretization of (2)
@@
Figure 1: The mesh for finite difference schemes. Cell centers are labeled as integers, and the interfaces or cell boundaries are labeled using half indices: $x_{i+\frac{1}{2}}$ is between cells $I_{i}$ and $I_{i+1}$. When reconstructing the function at an interface, we can either reconstruct from the left (as indicated in blue) or from the right (as indicated in green).
We denote the undivided jump and average across the interface $x_{i+\frac{1}{2}}$ by
$\Delta(\cdot)_{i+\frac{1}{2}}=(\cdot)_{i+1}-(\cdot)_{i},\hskip 8.53581pt\overline{(\cdot)}_{i+\frac{1}{2}}=\frac{(\cdot)_{i+1}+(\cdot)_{i}}{2}.$
Additionally, we introduce the entropy potential $\Psi({\bm{u}}):={\bm{v}}({\bm{u}})^{\top}{\bm{f}}({\bm{u}})-q({\bm{u}})$, where ${\bm{v}}({\bm{u}})=\nabla_{\bm{u}}\eta({\bm{u}})$ is the vector of entropy variables. A sufficient condition for constructing entropy conservative fluxes is provided by the following theorem.
Theorem 1.
([53]) The second-order accurate numerical scheme (3) with the two-point numerical flux $\tilde{{\bm{f}}}_{i+\frac{1}{2}}=\tilde{{\bm{f}}}_{i+\frac{1}{2}}({\bm{u}}_{i},{\bm{u}}_{i+1})$ is entropy conservative if
$\left(\Delta{\bm{v}}_{i+\frac{1}{2}}\right)^{\top}\tilde{{\bm{f}}}_{i+\frac{1}{2}}=\Delta\Psi_{i+\frac{1}{2}}.$ (5)
Specifically, it satisfies the entropy condition with the consistent numerical entropy flux given by
$\tilde{q}_{i+\frac{1}{2}}=\left(\overline{{\bm{v}}}_{i+\frac{1}{2}}\right)^{\top}\tilde{{\bm{f}}}_{i+\frac{1}{2}}-\overline{\Psi}_{i+\frac{1}{2}}.$
The expression (5) in Theorem 1 provides a recipe to construct second-order entropy conservative fluxes. It leads to a unique flux construction for scalar conservation laws (given $\eta)$. For systems of conservation laws, it is possible to carefully construct numerical fluxes that satisfy (5) with additional desirable properties [22, 5, 55, 19]. To construct higher-order entropy conservative fluxes, we follow the approach in [30]. This involves computing suitable linear combinations of the second-order two-point fluxes
$\tilde{{\bm{f}}}_{i+\frac{1}{2}}^{2p}=\sum_{r=1}^{p}B_{r}^{p}\sum_{s=0}^{r-1}\tilde{{\bm{f}}}({\bm{u}}_{i-s},{\bm{u}}_{i-s+r}),$ (6)
which yields $2p$th-order accurate entropy conservative fluxes. The scalar coefficients $B_{r}^{p}$ can be obtained by solving the $p$ linear equations
$2\sum_{r=1}^{p}rB_{r}^{p}=1,\hskip 8.53581pt\sum_{i=1}^{p}i^{2s-1}B_{r}^{p}=0\hskip 8.53581pt(s=2,...,p).$
For example, the fourth-order entropy conservative flux is given by
$\tilde{{\bm{f}}}_{i+\frac{1}{2}}^{2p}=\frac{4}{3}\tilde{{\bm{f}}}({\bm{u}}_{i},{\bm{u}}_{i+1})-\frac{1}{6}\left(\tilde{{\bm{f}}}({\bm{u}}_{i-1},{\bm{u}}_{i+1})+\tilde{{\bm{f}}}({\bm{u}}_{i},{\bm{u}}_{i+2})\right).$ (7)
@@
${\bm{f}}_{i+\frac{1}{2}}=\tilde{{\bm{f}}}^{2p}_{i+\frac{1}{2}}-\frac{1}{2}\mathbf{D}_{i+\frac{1}{2}}\Delta{\bm{v}}_{i+\frac{1}{2}},$ (8)
where $\mathbf{D}_{i+\frac{1}{2}}\succeq 0$ (a positive semi-definite matrix) is evaluated at some suitable averaged states. The numerical flux (8) satisfies the following result.
Lemma 1.
([14]) The numerical scheme (3) with numerical flux (8) satisfies the discrete entropy inequality
@@
where the numerical entropy flux is of the form
$q_{i+\frac{1}{2}}=\tilde{q}_{i+\frac{1}{2}}-\frac{1}{2}\left(\overline{{\bm{v}}}_{i+\frac{1}{2}}\right)^{\top}\mathbf{D}_{i+\frac{1}{2}}\Delta{\bm{v}}_{i+\frac{1}{2}}.$ (10)
While any positive semi-definite matrix ensures entropy stability in the sense of (9), we choose the form $\mathbf{D}_{i+\frac{1}{2}}=\mathbf{R}_{i+\frac{1}{2}}\Lambda_{i+\frac{1}{2}}\mathbf{R}^{\top}_{i+\frac{1}{2}}$. Here, $\mathbf{R}$ is a matrix consisting of the right eigenvectors of the flux Jacobian and $\mathbf{\Lambda}$ is a nonnegative diagonal matrix that depends on the eigenvalues of the flux Jacobian. Specifically, we choose the Roe-type diffusion matrix with $\mathbf{\Lambda}=\operatorname{diag}{(|\lambda^{1}|,...,|\lambda^{d}|)}$. See [53] for other choices for the diffusion matrix.
@@
We consider the interface at $x_{i+\frac{1}{2}}$ between the cells $I_{i}$ and $I_{i+1}$, as shown in Figure 1. We reconstruct from the left and right of this interface. We define the (locally) scaled entropy variables ${\bm{z}}=\mathbf{R}^{\top}_{i+\frac{1}{2}}{\bm{v}}$ corresponding to this interface. The flux in (8) can thus be expressed as
${\bm{f}}_{i+\frac{1}{2}}=\tilde{{\bm{f}}}^{2p}_{i+\frac{1}{2}}-\frac{1}{2}\mathbf{R}_{i+\frac{1}{2}}\Lambda_{i+\frac{1}{2}}\Delta{\bm{z}}_{i+\frac{1}{2}}.$ (11)
@@
([14]) For each interface $x_{i+\frac{1}{2}}$, if the reconstruction satisfies the sign property
@@
then the scheme with the numerical flux
${\bm{f}}_{i+\frac{1}{2}}=\tilde{{\bm{f}}}^{2p}_{i+\frac{1}{2}}-\frac{1}{2}\mathbf{R}_{i+\frac{1}{2}}\Lambda_{i+\frac{1}{2}}\llbracket{\bm{z}}\rrbracket_{i+\frac{1}{2}},$ (13)
is entropy stable.
High-order entropy stable schemes described in Lemma 2 are called TeCNO schemes, which rely on reconstructions satisfying the sign property (12). However, only a handful of reconstructions are known to satisfy the sign property. In [15], it was proven that ENO reconstructions satisfy this critical condition. In [16], third-order WENO schemes (known as SP-WENO) were designed to possess the sign property, which serves as the starting point for the novel reconstruction proposed in the present work. We discuss the SP-WENO formulation in Section 3.1.

### v2

# references.bib

@@ Abstract
Entropy conditions play a crucial role in the extraction of a physically relevant solution for systems of conservation laws, thus motivating the construction of entropy stable schemes that satisfy a discrete analogue of such conditions. TeCNO schemes (Fjordholm et al. 2012) [fjordholm2012arbitrarily] form a class of arbitrary high-order entropy stable finite difference solvers, which require specialized reconstruction algorithms satisfying the sign property at each cell interface. Third-order weighted essentially non-oscillatory (WENO) schemes called SP-WENO (Fjordholm and Ray, 2016) [fjordholm2016sign] and SP-WENOc (Ray, 2018) [ray2018third] have been designed to satisfy the sign property. However, these WENO algorithms can perform poorly near shocks, with the numerical solutions exhibiting large spurious oscillations. In the present work, we propose a variant of the SP-WENO, termed as Deep Sign-Preserving WENO (DSP-WENO), where a neural network is trained to learn the WENO weighting strategy. The sign property and third-order accuracy are strongly imposed in the algorithm, which constrains the WENO weight selection region to a convex polygon. Thereafter, a neural network is trained to select the WENO weights from this convex region with the goal of improving the shock-capturing capabilities without sacrificing the rate of convergence in smooth regions. The proposed synergistic approach retains the mathematical framework of the TeCNO scheme while integrating deep learning to remedy the computational issues of the WENO-based reconstruction. We present several numerical experiments to demonstrate the significant improvement with DSP-WENO over the existing variants of WENO satisfying the sign property.
@@ 1 Introduction
Hyperbolic systems of conservation laws dictate the behavior of quantities that are conserved in time as the system evolves. These systems of partial differential equations (PDEs) are ubiquitous across many disciplines. A generic system of conservation laws in one spatial dimension can be mathematically expressed as
@@
However, it is well-known that solutions to nonlinear conservation laws can develop discontinuities in finite time even with smooth initial conditions. Thus, solutions to (1) must be understood in the weak (distributional) sense. Moreover, since weak solutions are not unique in general, additional constraints in terms of the entropy conditions must be imposed such that a physically relevant weak solution is extracted. Assume that for (1), $\eta({\bm{u}})$ is a convex entropy function and $q({\bm{u}})$ is the entropy flux function satisfying the compatibility condition $\nabla_{{\bm{u}}}q({\bm{u}})^{\top}=\nabla_{{\bm{u}}}\eta({\bm{u}})^{\top}\nabla_{\bm{u}}{\bm{f}}({\bm{u}})$. The solution ${\bm{u}}$ of (1) is said to be an entropy solution if it satisfies the following inequality for all admissible entropy pairs $(\eta({\bm{u}}),q({\bm{u}}))$ associated with (1):
@@
which is understood in the weak sense. The entropy is conserved for smooth solutions, and thus (2) is taken to be an equality, i.e., $\eta$ is the solution of a conservation law. In contrast, entropy is dissipated near shocks, thus the inequality in (2) is strict near such discontinuities. Scalar conservation laws have unique entropy solutions [Kruzkov], but the uniqueness is not guaranteed for general systems of conservation laws [deLellis2010, Chiodaroli]. However, the entropy conditions provide the only non-linear estimates currently available for generic systems of conservation laws [dafermos2005hyperbolic], and are thus essential. In practice, the entropy condition (2) is considered with respect to a specific choice of $(\eta({\bm{u}}),q({\bm{u}}))$ for a general system (1).
@@
It is meaningful to develop entropy stable schemes, i.e., schemes satisfying a discrete version of (2), to solve systems of conservation laws. Further, it is essential for such methods to be high-order accurate while being capable of capturing discontinuities without spurious oscillations. In finite difference methods, the computational domain is partitioned into cells, on which a discrete version of the conservation law is formulated. Point values of the solution in the cells are evolved in time using time integration techniques, such as strong stability preserving Runge-Kutta methods [gottlieb2001strong]. In [tadmor2003entropy], a novel approach to constructing entropy stable finite difference schemes was proposed, which comprises two steps: (1) beginning with a second-order entropy conservative flux such that entropy is conserved locally and (2) adding an artificial dissipation term to ensure entropy stability.
@@
TeCNO schemes [fjordholm2012arbitrarily] are a popular class of arbitrary high-order entropy stable finite difference schemes which build on the formulation considered in [tadmor2003entropy]. These schemes consist of augmenting high-order entropy conservative fluxes with high-order numerical diffusion, which rely on specialized polynomial reconstruction algorithms. Crucially, these reconstructions must satisfy a sign property (see Lemma 1) at each cell interface so that entropy stability is ensured. Although any reconstruction method with the sign property can be used to formulate entropy stable TeCNO schemes, only a handful of such algorithms are currently available [fjordholm2012arbitrarily, cheng2016third, cheng2019fourth, fjordholm2016sign, ray2018third]. The essentially non-oscillatory (ENO) reconstruction method, which adaptively chooses the smoothest stencil for reconstruction, satisfies the sign property as demonstrated in [fjordholm2013eno]. To overcome some of the computational challenges encountered when using ENO, third-order weighted ENO (WENO) reconstructions were proposed in [fjordholm2016sign, ray2018third] which guarantee the sign property. However, these WENO schemes (termed as SP-WENO), while ensuring entropy stability in the TeCNO framework, are susceptible to spurious oscillations near discontinuities. We will highlight this issue in greater detail in the present work. We also mention here an alternate strategy [pandey2023sign] based on TeCNO schemes, where the sign property of reconstructions is not required. Instead, a specialized diffusion operator is constructed to ensure entropy stability, which can be used with any high-order reconstruction algorithm.
@@
The last few years have witnessed a surge in the use of deep learning-based strategies to solve problems in scientific computing. Key examples includes building surrogates to solve PDEs [raissi2019, pinns_review, deeponet, li2020fourier, varmion], learning parameter-to-observable maps [lye2020, lye2021iterative, huang2024operator], solving Bayesian inference problems [goh2022, patel2022solution, pigans, dasgupta2023conditional, raycgan23], and learning closure models [singh2016, beck2019, maulik2020]. Among these is the class of deep learning (DL) approaches which aim to enhance existing numerical solvers, instead of replacing them. The philosophy here is to first identify computational bottlenecks in the solver, and then use domain knowledge to train specialized neural networks to replace the bottleneck, keeping the rest of the solver intact. Thus, this synergistic approach leverages the “best of both worlds”, leading to a better numerical solver aided by DL. For instance, in [ray2018artificial, ray2019detecting], deep-learned (universal) troubled-cell indicators were designed for discontinuous Galerkin (DG) schemes solving conservation laws, to identify (classify) elements containing discontinuities. Similar DL-based shock capturing strategies have also been explored in [discacciati2020controlling, beck2020, zeifang2021, schwander2021, bruno2022, veiga2023, sun2020convolution, yu2022multi].
@@
DL techniques have also been used to learn ENO/WENO-type reconstructions. In [DeRyck2022], it was shown that ENO (and some of its variants) are equivalent to deep ReLU neural networks. A six-point ENO-type reconstruction was proposed in [li2022six] where the stencil selection is performed using a neural network. However, when used to solve conservation laws, some numerical experiments were accompanied by a degeneracy in the expected order of accuracy. In [stevens2020], a neural network was trained to develop an optimal variant of the classical WENO-JS [jiangshu96]. Despite leading to sharper shock profiles in numerical solutions, the expected accuracy was not always observed with smooth solution. A DL-based fifth-order WENO called WENO-DS was proposed in [kossaczka2021enhanced] with improved shock-capturing capabilities while retaining the formal order of accuracy. However WENO-DS is not model agnostic, and needs to be retrained if the conservation law is changed. An extension to the 2D Euler equations was proposed in [kossaczka2024deep].
@@
In the present work, we are interested in using DL to recover reconstruction algorithms constrained to satisfy useful physical properties. In particular, we construct a variant of SP-WENO, which we call DSP-WENO, where a neural network predicts the weights associated with the reconstruction, while ensuring that the sign property (along with other crucial constraints) are satisfied. These constraints are strongly imposed which results in a convex polygonal region for the WENO weight selection. Then, a neural network is trained on suitably generated training samples to learn the weight selection algorithm of the DSP-WENO for appropriate behavior near discontinuities and smooth regions. The proposed DSP-WENO is constructed to be third-order accurate, which is also demonstrated numerically. When used in the TeCNO framework, DSP-WENO overcomes the computational issues faced by ENO and the existing SP-WENO strategies. We re-iterate that the DSP-WENO only replaces the reconstruction needed in the high-order diffusion term, keeping the rest of the TeCNO framework intact. Thus the proposed method can be seen as a deep learning-based enhancement of an existing numerical algorithm. We remark here that a single DSP-WENO network is trained (offline) to be used with any conservation laws, i.e., the proposed algorithm is agnostic to the specific PDE model being solved.
@@ 2 Finite Difference Schemes/Entropy Stable Schemes
We consider a one-dimensional finite difference formulation for ease of discussion, which can easily be extended to higher dimensions. We partition the spatial domain $\Omega=[a,b]$ into $N$ disjoint cells $I_{i}=[x_{i-\frac{1}{2}},x_{i+\frac{1}{2}})$ of uniform length $h=(b-a)/N$ with cell center $x_{i}$ where
$x_{i+\frac{1}{2}}=a+ih,\quad\forall\ 0\leq i\leq N\ \ \ \text{and}\ \ \ x_{i}=\frac{x_{i-\frac{1}{2}}+x_{i+\frac{1}{2}}}{2}\quad\forall\ 1\leq i\leq N.$
Keeping time continuous, the semi-discrete finite difference scheme for (1) is expressed as
@@
Here, ${\bm{u}}_{i}(t)$ approximates the point values of the solution to (1) at cell center $x_{i}$ and time $t$, while ${\bm{f}}_{i+\frac{1}{2}}$ is a consistent, conservative numerical approximation of the flux ${\bm{f}}$ at the cell interface $x_{i+\frac{1}{2}}$. We seek entropy stable schemes to approximate (1) which satisfy a discrete version of the entropy condition (2). As described in [tadmor2003entropy], we begin by constructing an entropy conservative scheme which satisfies the discrete entropy equality
@@
We denote the undivided jump and average across the interface $x_{i+\frac{1}{2}}$ by $\Delta(\cdot)_{i+\frac{1}{2}}=(\cdot)_{i+1}-(\cdot)_{i}$ and ()¯$\overline{(\cdot)}_{i+\frac{1}{2}}=\big{(}(\cdot)_{i+1}+(\cdot)_{i}\big{)}/2$, respectively.
Additionally, we introduce the entropy potential $\Psi({\bm{u}}):={\bm{v}}({\bm{u}})^{\top}{\bm{f}}({\bm{u}})-q({\bm{u}})$, where ${\bm{v}}({\bm{u}})=\nabla_{\bm{u}}\eta({\bm{u}})$ is the vector of entropy variables. A sufficient condition [tadmor2003entropy] for the two-point flux $\tilde{{\bm{f}}}_{i+\frac{1}{2}}=\tilde{{\bm{f}}}_{i+\frac{1}{2}}({\bm{u}}_{i},{\bm{u}}_{i+1})$ to be entropy conservative is given by
$\big{(}\Delta{\bm{v}}_{i+\frac{1}{2}}\big{)}^{\top}\tilde{{\bm{f}}}_{i+\frac{1}{2}}=\Delta\Psi_{i+\frac{1}{2}}.$ (5)
The expression (5) provides a recipe to construct second-order entropy conservative fluxes, and leads to a unique numerical flux for scalar conservation laws (given $\eta)$. For systems of conservation laws, it is possible to carefully construct numerical fluxes that satisfy (5) with additional desirable properties [ismail2009, chandrashekar2013kinetic, winter2016, gouasmi2019]. Arbitrary high-order entropy conservative fluxes $\tilde{{\bm{f}}}^{p}$ (of order $p$) can be constructed using linear combinations of second-order two-point entropy conservative fluxes [lefloch2002fully]. For example, the fourth-order $(p=4)$ entropy conservative flux is given by
$\tilde{{\bm{f}}}_{i+\frac{1}{2}}^{4}=\frac{4}{3}\tilde{{\bm{f}}}({\bm{u}}_{i},{\bm{u}}_{i+1})-\frac{1}{6}\left(\tilde{{\bm{f}}}({\bm{u}}_{i-1},{\bm{u}}_{i+1})+\tilde{{\bm{f}}}({\bm{u}}_{i},{\bm{u}}_{i+2})\right).$ (6)
@@
${\bm{f}}_{i+\frac{1}{2}}=\tilde{{\bm{f}}}^{p}_{i+\frac{1}{2}}-\frac{1}{2}\mathbf{D}_{i+\frac{1}{2}}\Delta{\bm{v}}_{i+\frac{1}{2}},$ (7)
where $\mathbf{D}_{i+\frac{1}{2}}\succeq 0$ (a positive semi-definite matrix) is evaluated at some suitable averaged states. The numerical flux (7) leads to the satisfaction of the following discrete entropy inequality [fjordholm2012arbitrarily]
@@
where $q_{i+\frac{1}{2}}$ is a consistent numerical entropy flux. While any positive semi-definite matrix ensures entropy stability in the sense of (8), we choose the form $\mathbf{D}_{i+\frac{1}{2}}=\mathbf{R}_{i+\frac{1}{2}}\Lambda_{i+\frac{1}{2}}\mathbf{R}^{\top}_{i+\frac{1}{2}}$. Here, $\mathbf{R}$ is a matrix consisting of the right eigenvectors of the flux Jacobian and $\mathbf{\Lambda}$ is a nonnegative diagonal matrix that depends on the eigenvalues of the flux Jacobian. Specifically, we choose the Roe-type diffusion matrix with $\mathbf{\Lambda}=\operatorname{diag}{(|\lambda^{1}|,...,|\lambda^{d}|)}$. See [tadmor2003entropy] for other choices for the diffusion matrix.
@@
We consider the interface at $x_{i+\frac{1}{2}}$ between the cells $I_{i}$ and $I_{i+1}$ and reconstruct from the left and right of this interface. We define the (locally) scaled entropy variables ${\bm{z}}=\mathbf{R}^{\top}_{i+\frac{1}{2}}{\bm{v}}$ corresponding to this interface. The flux (7) can thus be expressed as
${\bm{f}}_{i+\frac{1}{2}}=\tilde{{\bm{f}}}^{p}_{i+\frac{1}{2}}-\frac{1}{2}\mathbf{R}_{i+\frac{1}{2}}\Lambda_{i+\frac{1}{2}}\Delta{\bm{z}}_{i+\frac{1}{2}}.$ (9)
@@
([fjordholm2012arbitrarily]) For each interface $x_{i+\frac{1}{2}}$, if the reconstruction satisfies the sign property
@@
then the scheme with the following numerical flux is entropy stable
${\bm{f}}_{i+\frac{1}{2}}=\tilde{{\bm{f}}}^{p}_{i+\frac{1}{2}}-\frac{1}{2}\mathbf{R}_{i+\frac{1}{2}}\Lambda_{i+\frac{1}{2}}\llbracket{\bm{z}}\rrbracket_{i+\frac{1}{2}}.$ (11)
High-order entropy stable schemes described in Lemma 1 are called TeCNO schemes, which rely on reconstructions satisfying the sign property (10). However, only a handful of reconstructions are known to satisfy the sign property. In [fjordholm2013eno], it was proven that ENO reconstructions satisfy this critical condition. In [fjordholm2016sign], third-order WENO schemes (known as SP-WENO) were designed to possess the sign property, which serves as the starting point for the novel reconstruction proposed in the present work. We discuss the SP-WENO formulation in Section 3.1.